OpenStudy (anonymous):

Put the equation below in the form ax2 + bx + c = 0. Enter exponents using the caret ( ^ ). For example, you would enter 4x2 as 4x^2. x2 + 2x + 9 = 2

OpenStudy (anonymous):

i need help.

ganeshie8 (ganeshie8):

x2 + 2x + 9 = 2

ganeshie8 (ganeshie8):

you need to convert it to ax2 + bx + c = 0

ganeshie8 (ganeshie8):

right ?

OpenStudy (anonymous):

yes.

ganeshie8 (ganeshie8):

first make the right side 0 subtract 2 both sides

ganeshie8 (ganeshie8):

x2 + 2x + 9 = 2 x2 + 2x + 9-2 = 2-2 x2 + 2x + 7 = 0

OpenStudy (anonymous):

is thats the answer.

ganeshie8 (ganeshie8):

u think.. x2 has no co-efficent ?... is that throwing u off ?

OpenStudy (anonymous):

so then it would 2x^2

ganeshie8 (ganeshie8):

it is in form ax2 + bx + c =0 so that must be d answer

ganeshie8 (ganeshie8):

nope x^2 \(\ne \) 2x^2

ganeshie8 (ganeshie8):

x^2 = 1x^2

ganeshie8 (ganeshie8):

when you dont see any co-effecient, you can think of it as : x^2 = 1x^2 so co-effecient is 1

OpenStudy (anonymous):

yea so its x2 + 2x + 7 = 0

ganeshie8 (ganeshie8):

yes. its same as, 1x2 + 2x + 7 = 0 here : a =1 b =2 c = 7

OpenStudy (anonymous):

oh'k ill plug it in now.

ganeshie8 (ganeshie8):

good luck :) careful... it asks you to enter exponents using ^ sign..

OpenStudy (anonymous):

oh i got it. thank you.

ganeshie8 (ganeshie8):

np..

OpenStudy (anonymous):

will you help me with the next one?

ganeshie8 (ganeshie8):

sure post it :)

OpenStudy (anonymous):

x2 - 2 = -2x2 + 5x

ganeshie8 (ganeshie8):

add 2x2 both sides

ganeshie8 (ganeshie8):

lets do things that make right side = 0

OpenStudy (anonymous):

2x^2-x^2=5x

OpenStudy (anonymous):

?

ganeshie8 (ganeshie8):

where did -2 go... we will go step by step... so that it becoems clear

OpenStudy (anonymous):

oh'k.

ganeshie8 (ganeshie8):

x2 - 2 = -2x2 + 5x add 2x2 both sides x2 - 2 + 2x2 = -2x2 + 5x + 2x2 x2 + 2x2 - 2 = 5x 3x2 - 2 = 5x subtract 5x both sides 3x2 - 2 - 5x = 5x - 5x 3x2 - 5x - 2 = 0

OpenStudy (anonymous):

yay! i got it right.

ganeshie8 (ganeshie8):

:) glad to hear.. .

OpenStudy (anonymous):

3x2 + 27x + 60 = 0?